# Geometric Reasoning in the Embedding Space   Jan Hůla1,2, David Mojžíšek2, Jiří Janeček2, David Herel1, Mikoláš Janota1

### v1

@@ 4.1 Comparison of the Two Architectures
The experiments on the synthetically generated data show that the GNN performs significantly better than the Transformer, as one could expect. In 4.4, we show that we were able to train the GNN on grid sizes up to $80\times 80$ points to a validation accuracy larger than $90\,\%$. In comparison, the Transformer achieved accuracy of $90\,\%$ only if we train it on a grid size $10\times 10$ and limit the number of constraints to $2$. If we train the model on a more complex setting with the grid size $20\times 20$ and up to $6$ constraints, it reaches an accuracy of approximately $30\,\%$.
@@
For completeness, we also trained the Transformer model to find the assignments to variables using Chain-of-Thought (CoT) Wei et al. (2022) by imitating a log from a simple solver. The solver resolves the constraints one by one in the topological order of the DAG of constraints mentioned in Section 3.1. Using this way of training, the Transformer learns to predict the positions of variables within $20\times 20$ with approximately 50% point accuracy. We did not explore this direction further as reasoning with a CoT is orthogonal to reasoning in the embedding space on which we focus in this work. More details about the CoT experiment can be found in Appendix F.
@@
Given the superior results achieved by the GNN, we conduct the main analysis of the embeddings on the GNN. Appendix E contains similar analysis and visualization for the Transformer (see Figure 7).
@@
We also note that the model accuracy can be further improved by leveraging multiple prediction attempts since incorrect predictions often fall close to their true positions (as shown in Appendix D) and according to preliminary tests, sampling multiple different initial embeddings increases the chance of predicting the correct assignment.
@@ 4.2 Visualizing the Embeddings of Individual Points
When visualizing low-dimensional projection of the static embeddings corresponding to individual points, we found that they organize themselves into a 2D grid they represent. In Figure 1, we show how this organization emerges during training and how it is connected to the precision of the prediction (three dimensional projection is depicted in Appendix C).
Figure 1: Evolution of static embeddings throughout training. Lines and colors represent the connectivity of points within the 30$\times$30 grid in one direction. UMAP McInnes et al. (2018) is used as a projection method from the original 96 dimensions of the embedding space. We found that UMAP works best for larger grids; however, simpler PCA projections worked better for smaller grids (15$\times$15 and less points).
@@
4.2.1 Solution Process
In order to find the solution, the GNN model iteratively moves dynamic embeddings in a high-dimensional space. During the forward pass, the same update rule is applied over and over again. Hence, it is possible to extract the embeddings in each step of the model to get a better understanding of the underlying process.
@@ 4.2.1 Solution Process
Note that in this example, the GNN first finds a close approximation to the hidden configuration and then refines it. Here, the squares are first “approximated” by quadrilaterals which then converge to exact squares. It can also be observed that one square constraint is approximately satisfied earlier than the other which reflects the fact that the other constraint can be resolved only after the first constraint is resolved (as explained in the next section).
Figure 2: Visualization of the solution process. The red points A, B, C, D, E, F (C = E) are known, and the blue points G, H, I, J, K, L need to be predicted. There are two translation constraints: T (F, D, C=E, G), T (C=E, G, A, H) and two square constraints: S (H, I, G, J), S (B, J, L, K). Translations are marked by a dotted line and squares by a solid line. The network is trained to predict the result in $15$ iterations, of which initial state and results after iterations $3,5,7,9,11,12,13$ were chosen for illustration. The network gradually improves the result over the iterations: the first translation with only one unknown point G is solved, followed by finding the point H of the second translation. After translations, both squares are solved.
@@ 4.3 Analysis of Incorrectly Classified Points
(a) Relative incorrect point distribution.
(b) Total point distribution.
Figure 3: Predicted points classified based on the number of resolving steps (on the constraint level) that are
necessary to determine the point positions.
To understand the failure modes of the GNN, we analyze predictions of approximatelly 10k instances of our CSP. These instances contain around 29k points to be classified. As explained in Section 3.1, the constraints of each instance could be organized into a DAG by a dependency given by the variables. A constraint A depends on a constraint B if the determining variables of A are the dependent variables of B. To resolve a given variable, we can start with the constraint which contains it and recursively resolve all the constraints on which this constraint depends. Once all these constraints are resolved, we can also resolve the constraint in which this variable appears, and thus we get its value. The constraints which need to be resolved in order to resolve a given variable will constitute a sugraph of the DAG and the number of constraints in this subgraph determines the number of steps which are needed to resolve the given variable. If the variable is contained in multiple constraints, then we can consider the one which requires the least number of resolving steps.
We hypothesized that the incorrect predictions will be correlated with the number of resolving steps required for a given variable. The barplot in Figure 3a confirms this hypothesis. It shows that the relative failure rate (incorrect predictions vs. all predictions for a particular point) depends almost monotonically on the number of resolving steps required. Monotonicity of this dependency is violated only in the case of 7 resolving steps which could be attributed to the insufficient number of tested points to robustly estimate the failure rate. The absolute numbers of predicted points in the generated problems are shown in Figure 3b. We also mention that the incorrectly classified points are very often classified to a point lying very close the correct point (see Appendix D).
4.4 Scaling the Size of the Grid
The problem generator mentioned in 3.1 produces problems which have on average around four constraints. Both types of constraints ($S$ and $T$) are sampled with equal probability. Therefore, we can assume that an average problem has two constraints of type $S$ (which is determined by $2$ points) and two constraints of type $T$ (which is determined by $3$ points). If we denote the number of points on the side of the grid by $n$, then we can estimate the number of unique problems in the grid of size $n\times n$ to be $n^{20}$. There are $n^{2}$ possible point positions and we independently sample 6 points for two constraints of type T and 4 points for two constraints of type S, together yielding $(n^{2})^{10}$ possibilities each of which determines one instance. This number should be viewed as an upper bound because it ignores the fact the constraints can share variables.
To test the dependence of the validation accuracy on the grid size and number of training samples, we use the following values of $n$: $10$, $20$, $30$, $40$, $50$, $60$, $70$, $80$. This results in the following number [^4] After rounding the exponent to the nearest integer. of possible problems for each $n$, respectively: $10^{20}$, $10^{26}$, $10^{29}$, $10^{32}$, $10^{34}$, $10^{35}$, $10^{37}$, $10^{38}$.
The sizes of the training set for each grid size are in the range from $5$k to $800$k. The relationship between the validation accuracy, the grid size and the size of the training set can be seen in Figure 4a. In Figure 4b, we plot the relationship between the grid size and sizes of the training set for which the validation accuracy exceeded $90\,\%$. As can be seen, in this range of grid sizes, the sample complexity grows much more slowly than the number of possible examples.
4.5 Impact of Constraint Types
We also analyzed how model performance changes when incorporating additional geometric constraint types while maintaining consistent training conditions on a $20\times 20$ grid and a fixed training set size of 70k examples (Table 1). Adding each new constraint type (M, R) to the base set (S, T) led to substantial decreases in both point-wise and complete-figure accuracy. The observed decrease in accuracy appears steeper than would be expected from linear scaling with constraint count, suggesting increased problem complexity. Given our findings on scaling laws in Section 4.4, it is likely that higher accuracy could be achieved with larger training datasets.
(a) Accuracy vs number of training samples for
different grid sizes.
(b) Grid size vs sample size required for accuracy
$>90\,\%$.
Figure 4: Scaling laws for different sizes of the 2D grid. For a grid of size $n$, there are $n^{2}$ points.
@@ 5 Limitations and Discussion
We note that we study a much simpler setup than was studied in AlphaGeometry. Predicting positions of points satisfying a set of geometric constraints is straightforward in comparison with predicting useful auxiliary points. Also, in our case, the unknown points need to lie on a discrete 2D grid so that we can train the model with a classification loss. The simplified setup turned out to be sufficient to study how NNs can reason about geometric constraints. Training the model for predicting auxiliary points requires a large computational budget and, moreover, the authors did not release the training dataset, which is also expensive to create. We also mention that we use only four types of constraints, but the whole setup could be easily extended to different sets of constraints by adding more update functions in Equation (2). These simplifications enabled us to provide partial understanding of the process by which the tested models are able to solve individual problems and to demonstrate the benefits of using a GNN instead of a Transformer.
@@
Compared to the GNN which Hůla et al. (2024) used for Boolean satisfiability, our GNN can deal with fewer constraints and points. In their case, the GNN was able to handle hundreds of constraints, whereas our GNN struggles with instances that contain more than 10 constraints. This could be caused by the fact that the domain of our constraints is much larger, but it could also be the case that different architecture of the GNN (with different update functions) could scale to larger problems. We leave this exploration for future work.
@@
It is also possible that a different training procedure might result in significantly easier training. The inference process depicted in Figure 2 and the findings of Hůla et al. (2024) suggest that the GNN could implicitly optimize an unknown objective during the iterative forward pass. Finding an expression for this objective is another possible direction for future work. It could be used to train the network in an unsupervised way which should be much easier that training it with the classification loss. Another obvious direction is to train the model as a diffusion model which “denoises” randomly assigned points to points of the solution.
@@
Lastly, we mention that our experimental setup can be extended to more complex CSPs which could contain temporal relations and relations between various entities. We believe that such CSPs could be very useful for studying how NNs can generalize to unseen situations Abbe et al. (2023) and how they can discover models of the world without grounding Wong et al. (2023). Studying geometric CSPs has the advantage that the domain has an obvious geometric interpretation which is visible in the embedding space.
6 Conclusion
@@
We have shown that GNNs as well as Transofmers can learn to solve geometric CSPs and provided several insights into the process by which they find the solution. The visualizations show that when processing the problem, models form the hidden spatial configurations described by the problem in the embedding space. During training, the static embeddings of individual points in the 2D grid organize themselves within a 2D subspace and reflect the neighborhood structure of the grid. We showed that the occurence of errors depends on the number of steps required to resolve a given variable and the number of types of constraints present within the CSP language. Lastly, we showed that GNNs are much easier to train and can be scaled to significantly larger grids than Transformers.
Appendix A Appendix
Appendix B Model and Training Hyperparameters
The final hyperparameters for both model architectures and for their training can be seen in Tables 3 and 3.
Table 2: Model hyperparameters and training settings for the GNN.
Graph Neural Network
Hyper-parameter Value
Embedding dimension 96,256
Embedding update iterations 20
Batch size 256
Number of epochs 200
Optimizer AdamW
Initial learning rate $10^{-3}$
Warmup steps 5
Learning rate scheduler CosineAnnealingLR
Gradient clipping norm $0.6$
Weight Decay $10^{-6}$
Table 3: Model hyperparameters and training settings for the Transformer.
Transformer
Hyper-parameter Value
Number of layers 6

### v2

@@ 4.1 Comparison of the Two Architectures
The experiments on the synthetically generated data show that the GNN performs significantly better than the Transformer, as expected. In Appendix G, we show that we were able to train the GNN on grid sizes up to $80\times 80$ points to a validation accuracy larger than $90\,\%$. In comparison, the Transformer achieved accuracy of $90\,\%$ only if we train it on a grid size $10\times 10$ and limit the number of constraints to $2$. If we train the model on a more complex setting with the grid size $20\times 20$ and up to $6$ constraints, it reaches an accuracy of approximately $30\,\%$.
@@
For completeness, we also trained the Transformer model to find the assignments to variables using Chain-of-Thought (CoT) Wei et al. (2022) by imitating a log from a simple solver. The solver resolves the constraints one by one in the topological order of the DAG of constraints mentioned in Section 3.1. Using this way of training, the Transformer learns to predict the positions of variables within $20\times 20$ grid with approximately 50% point accuracy. We did not explore this direction further as reasoning with a CoT is orthogonal to reasoning in the embedding space on which we focus in this work. More details about the CoT experiment can be found in Appendix E.
@@
Given the substantially better results achieved by the GNN on this problem setting, we conduct the main analysis of the embeddings with the GNN. Appendix D contains similar analysis and visualization for the Transformer (see Figure 16).
@@
We also note that the model accuracy can be further improved by leveraging multiple prediction attempts since incorrect predictions often fall close to their true positions (as shown in Appendix C) and according to preliminary tests, sampling multiple different initial embeddings increases the chance of predicting the correct assignment.
4.2 GNN Performance Analysis
Given the GNN’s clear advantage over the Transformer for our CSPs representable by bipartite graphs, we now analyze the GNN’s behavior in detail across multiple dimensions. We examine how the model performs under distribution shift, visualize the internal representations it develops, and identify failure modes related to problem complexity. The analysis uses our best-performing GNN configuration with embedding dimension $128$, $15$ message-passing iterations, and training procedures detailed in Appendix A. This systematic evaluation reveals both the capabilities and limitations of our geometric constraint solving approach on discrete grids.
4.2.1 Dataset Design and Complexity Distribution
The training dataset contains problems with 111616 geometric constraints (mean: $4.0$), 333434 points (mean: $8.7$), and reasoning depths of 111212 (mean: $2.9$). The test dataset was specifically designed to be more challenging, containing problems with 882626 geometric constraints (mean: $14.2$), 12124747 points (mean: $23.8$), and reasoning depths of 331414 (mean: $7.0$). Figure 4 illustrates these distribution differences across key complexity metrics.
Our generator creates problems for specific grid sizes, and we focus on a $20\times 20$ grid as it provides a good balance. It is large enough to accommodate interesting geometric figures while remaining computationally tractable for detailed analysis, which required multiple re-runs. Our constraint set consists of four types: Square (S), Reflection (R), Midpoint (M), and Translation (T), with their relative distribution in the training dataset shown in Figure 3.
Figure 3: Distribution of constraint types in the training dataset. Square constraints are most frequent (41.7%), followed by Midpoint (26.7 %), Translation (18.8 %), and Reflection (12.9 %). The test dataset follows a very similar distribution, ensuring consistent constraint type representation across training and evaluation.
This choice of difficulty distributions allows us to study model behavior when problem complexity is increased beyond what model encountered during training. The high validation accuracy achieved on the training distribution ($99.5$ % point accuracy, $98.1$ % complete problem accuracy using $10$  % of training data as validation) demonstrates that the GNN can effectively solve geometric constraint problems when sufficient training data is available. We used $300000$ ($270000$ without validation) training examples. The size was determined through scaling analysis outlined in Appendix G, which shows the sample complexity requirements for different grid sizes.
The deliberately harder test distribution provides a challenging evaluation setting, where the model achieves lower baseline accuracy, creating opportunities to study failure modes, reasoning depth effects, and test-time scaling approaches.
Figure 4: Deliberate distribution shift between training and test sets across complexity metrics. The harder test distribution enables analysis of model behavior under increased problem complexity, failure mode identification, and evaluation of test-time scaling approaches.
4.2.2 Performance under Distribution Shift
Table 1 presents comprehensive performance results on our datasets. The model maintains good performance even under distribution shift, achieving $96.07$ % point accuracy and $76.74$ % complete accuracy on the harder test set with standard inference ($15$ iterations, $1$ resample).
The results demonstrate clear benefits from test-time scaling approaches. Increasing inference iterations from $15$ to $23$ improves complete accuracy from $76.74$  % to $93.62$  %. Using $10$ resamples with different random initializations further boosts performance to $95.37$ % complete accuracy. We determined the optimal iteration count of $23$ through analysis shown in Figure 18 which also reveals a slight performance increase with multiple resamples.
The “Best” configuration could be explained as an oracle that knows the optimal number of iterations for each individual problem to achieve the most solved points in the fewest iterations (with a maximum of $50$ iterations). This approach achieves the complete accuracy of $96.14$  % with single resampling and $97.46$ % with $10$ resamples. This model could in theory get very close to it performance on validation set.
For resampling strategy, we tested various counts but found the most significant improvement occurs when going from $1$ to $5$ resamples. We chose $10$ resamples as a conservative buffer, since additional resamples beyond this point show diminishing returns while increasing computational cost.
These findings suggest several important properties of the learned model. The improvement from additional iterations indicates that the GNN employs an iterative refinement process similar to continuous optimization methods. The benefits of resampling show that different random initializations of unknown variables can lead to different solution paths, with some being more effective for particular problem instances.
Model’s ability to generalize to problems with increased complexity beyond the training distribution, are consistent with previous work that applied a very similar GNN architecture to SAT problems Mojžíšek et al. (2025).
Table 1: Performance of the model. The first row reports validation accuracy (on a separate dataset, using 1 resample). The remaining rows show results on a harder test set drawn from a different distribution, across different numbers of inference iterations (15, 23, Best) and resample counts (1 and 10). We report point-wise accuracy and complete accuracy.
1 Resample 10 Resamples
Setting Point Acc. Complete Acc. Point Acc. Complete Acc.
Validation (15 iters) 99.55 % 98.93 %
Test (15 iters) 96.07 % 76.74 % 97.89 % 80.93 %
Test (23 iters) 98.51 % 93.62 % 99.02 % 95.37 %
Test (Best) 99.04 % 96.14 % 99.44 % 97.46 %
@@ 4.3 Visualizing the Embeddings of Individual Points
When visualizing low-dimensional projection of the static embeddings corresponding to individual points, we found that they organize themselves into a 2D grid they represent. In Figure 5, we show how this organization emerges during training and how it is connected to the precision of the prediction (three dimensional projection is depicted in Appendix B).
Figure 5: Emergence of spatial structure in static point embeddings during training on $\times20$ grid. Colors indicate distance from grid center, revealing how the model learns geometric relationships without explicit spatial supervision. UMAP projection of 128-dimensional embeddings clearly shows the progression from random initialization to organized 2D grid-like structure, while PCA (Appendix B) points to the underlying curved manifold geometry. This self-organization demonstrates that geometric inductive biases emerge naturally from constraint satisfaction training.
@@
Here we focus on point embeddings, for completeness we provide some results about constraint embeddings in Appendix I. We show that constraint embeddings encode constraint types, geometric properties, satisfaction status, and temporal information up to some extent.
4.4 Network Initialization with Grid Structure
We investigated whether providing the model with an initial understanding of grid structure accelerates training convergence. Rather than initializing the shared embedding and classifier weights randomly, we can initialize them to reflect the geometric structure of the $20\times 20$ grid.
To provide the model with an initial geometric inductive bias, we initialize the embedding matrix $\mathbf{W}\in\mathbb{R}^{N\times d}$ (used for both known point embeddings and the final classification layer) using a rotated grid construction. We begin by assigning each of the $N$ grid positions ($N=400$ for a $20\times 20$ grid) its exact 2D coordinates $(x_{i},y_{i})$ and construct a matrix $\mathbf{V}\in\mathbb{R}^{N\times d}$ where $\mathbf{V}_{i,0}=x_{i}$, $\mathbf{V}_{i,1}=y_{i}$, and all other dimensions are zero. We then generate a random orthogonal matrix $\mathbf{Q}\in\mathbb{R}^{d\times d}$ via QR decomposition of a matrix with entries sampled from a standard normal distribution, and define initial embedding matrix as:
$\mathbf{W}=\mathbf{V}\mathbf{Q}$
This rotation spreads the spatial coordinate information across all $d$ dimensions while avoiding alignment with any particular coordinate axis. Following this initialization, training proceeds identically to the random case, with the weights in $\mathbf{W}$ remaining fully trainable.
Figure 6 demonstrates that grid-structured initialization provides substantial training acceleration compared to random initialization. The model with grid initialization converges to high validation accuracy within the first few epochs, while random initialization requires significantly more training time. This suggests that providing geometric inductive bias through weight initialization helps the model quickly discover the spatial relationships it needs to solve geometric constraint problems.
The initialization method shares weights between the embedding layer (used for known points) and classifier head (used for predictions), ensuring consistent geometric representations throughout the network. This architectural choice proves beneficial as the model can leverage the geometric structure for both encoding known positions and predicting unknown ones.
Figure 6: Comparison between random initialization of shared embedding and classification layer weights (red) and grid-structured initialization (blue) in terms of validation loss (left) and point accuracy (right). Red dots mark when each model first achieves 90 % accuracy. Grid initialization reaches 90 % accuracy within 10 epochs while random initialization requires almost 50 epochs. After 100 epochs their performance equalizes.
4.5 Solution Process
In order to find the solution, the GNN model iteratively moves embeddings of unknown points in a high-dimensional space. During the forward pass, the same update rule is applied over and over again. Hence, it is possible to extract the embeddings in each message passing iteration to get a better understanding of the underlying process. More similar examples are provided in Appendix H.
@@ 4.5 Solution Process
Note that in this example, the GNN first finds a close approximation to the hidden configuration and then refines it. Here, the squares are first “approximated” by quadrilaterals which then converge to exact squares. It can also be observed that one square constraint is approximately satisfied earlier than the other which reflects the fact that the other constraint can be resolved only after the first constraint is resolved (as explained in the next section). For an extended discussion and visualization of the reasoning dynamics as revealed by UMAP, see Appendix J.
Figure 7: Visualization of the solution process. The red points A, B, C, D, E, F (C = E) are known, and the blue points G, H, I, J, K, L need to be predicted. There are two translation constraints: T (F, D, C=E, G), T (C=E, G, A, H) and two square constraints: S (H, I, G, J), S (B, J, L, K). Translations are marked by a dotted line and squares by a solid line. The network is trained to predict the result in $15$ iterations, of which initial state and results after iterations $3,5,7,9,11,12,13$ were chosen for illustration. The network gradually improves the result over the iterations: the first translation with only one unknown point G is solved, followed by finding the point H of the second translation. After translations, both squares are solved. Note this visualization was created with model operating on 30x30 grid.
@@ 4.6 Analysis of Incorrectly Classified Points
Figure 8: Failure analysis on test dataset (15 iterations, 1 resample). Left: individual point prediction accuracy degrades with reasoning depth, showing points requiring deeper dependency chains have higher failure rates. Right: complete problem success decreases with constraint count. Numbers above bars show absolute sample sizes for each category.
To understand the failure modes of the GNN, we analyze predictions on our $30,000$ test instances using the basic model setup ($15$ iterations, $1$ resample) from Table 1. We examine how prediction accuracy relates to problem complexity along multiple dimensions.
As explained in Section 3.1, constraints form a dependency structure where some must be resolved before others. For any given point, we can trace the dependency chain of constraints that must be resolved to determine its position. We define point depth as the number of constraint resolution steps required for a specific point. Problems themselves vary in complexity through both the total number of constraints and the maximum depth of dependency chains within the problem.
Figure 8 (left) shows that prediction accuracy correlates negatively with point depth. Points requiring deeper reasoning chains have significantly higher failure rates. The relationship is nearly monotonic, with success rates declining from over $90\,\%$ for points at depth 2233 to below $20\,~\%$ for points requiring $10$ or more resolution steps. This pattern aligns with the iterative solution process observed in Section 4.5 where the model progressively constructs solutions by first resolving simpler constraints.
Figure 8 (right) demonstrates that problems with more constraints also show lower complete success rates. This reflects both the increased likelihood of errors accumulating across multiple predictions and the tendency for larger problems to contain deeper dependency structures.
These findings suggest that the GNN’s reasoning capability is fundamentally limited by the length of dependency chains it can reliably process. Importantly, when predictions do fail, the incorrectly classified points tend to be positioned close to their correct target locations, as evidenced by the distance histograms in Appendix C.
@@ 5 Limitations and Discussion
We note that we study a much simpler setup than was studied in AlphaGeometry. Predicting positions of points satisfying a set of geometric constraints is straightforward in comparison with predicting useful auxiliary points. Also, in our case, the unknown points need to lie on a discrete 2D grid so that we can train the model with a classification loss. The simplified setup turned out to be sufficient to study how NNs can reason about geometric constraints. Training the model for predicting auxiliary points requires a large computational budget and, moreover, the authors did not release the training dataset, which is also expensive to create. We also mention that we use only four types of constraints, but the whole setup could be easily extended to different sets of constraints by adding more update functions in Equation (1). These simplifications enabled us to provide partial understanding of the process by which the tested models are able to solve individual problems and to demonstrate the benefits of using a GNN instead of a Transformer. We note that variable permutation invariance is not considered in our constraint definitions, as a fixed variable ordering is required to ensure unique solutions.
@@
Compared to the GNN which Hula et al. Hůla et al. (2024) used for Boolean satisfiability, our GNN deals with fewer constraints and points. In their case, the GNN was able to handle hundreds of constraints, whereas our GNN deals with an order of magnitude fewer constrains and still needs more training samples than in the case of Boolean satisfiability. This could be caused by the fact that the domain of our constraints is much larger, but it could also be the case that different architecture of the GNN (with different update functions) could scale better to larger problems.
@@
It is also possible that a different training procedure might result in significantly easier training. The inference process depicted in Figure 7 and the findings of Hůla et al. (2024) suggest that the GNN could implicitly optimize an unknown objective during the iterative forward pass. Finding an expression for this objective could be interesting and used to train the network in an unsupervised way. This should be much easier than training it with the classification loss. Another obvious direction is to train the model as a diffusion model which “denoises” randomly assigned points to points of the solution.
@@
Lastly, we mention that our experimental setup can be extended to more complex CSPs which could contain temporal relations and relations between various entities. We believe that such CSPs could be very useful for studying how NNs can generalize to unseen situations Abbe et al. (2023) and how they can discover models of the world without grounding Wong et al. (2023). Studying geometric CSPs has the advantage that the domain has an obvious geometric interpretation which is visible also in the embedding space.
6 Future Work
While we provided initial insights into geometric reasoning in neural networks, several directions remain for deeper understanding. Future work should investigate the expressiveness requirements of different update mechanisms, building on our finding that LSTMs significantly outperform RNNs for constraint updates. This includes determining the minimal architectural complexity needed to solve geometric constraints and whether these mechanisms can recover explicit mathematical formulas underlying constraints rather than learning approximations. The iterative refinement process that we observed suggests connections to continuous optimization methods. Future work should formalize this relationship and investigate whether training procedures based on explicit optimization objectives could improve efficiency. Additionally, extending beyond our four constraint types while addressing variable permutation invariance would broaden insights. Moving from discrete grid classification to continuous coordinate regression would enable analysis of geometric reasoning in unrestricted spatial domains.
7 Conclusions
@@
We have shown that GNNs as well as Transformers can learn to solve geometric CSPs and provided several insights into the process by which they find the solution. The visualizations show that when processing the problem, models form the hidden spatial configurations described by the problem in the embedding space. During training, the static embeddings of individual points in the 2D grid organize themselves within a 2D subspace and reflect the neighborhood structure of the grid. We showed that the occurrence of errors depends on problem complexity, like the number of steps required to resolve all variables. The models solve constraints through an iterative refinement process and exhibit test-time scaling capabilities, where allocating additional computational resources during inference improves performance on harder (out of training distribution) problem instances. Lastly, we showed that GNNs are much easier to train and can be scaled to significantly larger grids than Transformers.
Author Contributions
Conceptualization, J.H. and D.M.; methodology, J.H. and D.M.; software, J.H., D.M., D.H. and M.J.; validation, D.M. and J.H.; formal analysis, D.M. and J.J.; investigation, D.M. and J.H.; resources, J.H. and M.J.; data curation, D.M.; writing—original draft preparation, D.M. and J.H.; writing—review and editing, J.J., D.M. and J.H.; visualization, D.M.; supervision, J.H. and M.J.; project administration, J.H. All authors have read and agreed to the published version of the manuscript.
Funding
This article has been produced with the financial support of the European Union under the REFRESH – Research Excellence For REgion Sustainability and High-tech Industries project number CZ.10.03.01/00/22_003/0000048 via the Operational Programme Just Transition.
Data Availability
The synthetic datasets and problem generator code supporting the findings of this study are available on request from the corresponding author. The codebase including model implementations and training procedures will be made publicly available after publication.
Acknowledgments
During the preparation of this manuscript, the authors used Claude (Anthropic’s AI assistant) for code development assistance and English language editing suggestions. All conceptualization, experimental design, analysis, and interpretation of results were conducted by the authors. The AI tool was used primarily for programming assistance and language refinement suggestions. The authors have reviewed and edited all output and take full responsibility for the content of this publication.
Conflicts of Interest
The authors declare no conflicts of interest.
Abbreviations
The following abbreviations are used in this manuscript:
NN Neural network GNN Graph neural network CSP Constraint satisfaction problem SDP Semidefinite programming LLM Large language model RNN Recurrent neural network M Midpoint constraint type R Reflection constraint type S Square constraint type T Translation constraint type DAG Directed acyclic graph LSTM Long short-term memory CoT Chain-of-thought PCA Principal component analysis UMAP Uniform manifold approximation and projection
Appendix A Training Procedures and Hyperparameters
This section details the training procedures and hyperparameter configurations used for the GNN models, which were the primary focus of experiments.
A.1 GNN Training Procedure
The GNN uses AdamW optimizer with cosine annealing learning rate scheduling. The scheduler operates in 15-epoch cycles where each cycle’s peak learning rate decreases by factor 0.9 from the previous cycle. Learning rate restarts cause temporary 0.5–6 % accuracy (larger in initial stages of training) drops but enable continued optimization, with models regaining and exceeding previous performance within few epochs.
Exponential Moving Average (EMA) weights with decay 0.99 provide stable evaluation metrics. EMA weights are used exclusively for validation and testing while training continues with primary parameters. Cross-entropy loss applies only to unknown points, with known points remaining fixed.
The model supports the training with both fixed and changing number of iterations (number of message passing repetitions is different for each batch). Changing iteration training samples from a diminishing probability distribution centered at 15 iterations ($\pm$10 range), where 15 iterations occur 40 % of the time, 16–17 iterations occur $\sim$15 % combined, while extreme values like 25 iterations occur <<1 %. This technique improves test-time scaling robustness. Iteration count does not affect trainable parameters since the same layers are applied recurrently.
Dropout (0.1) applies to both point and constraint embeddings during training.
A.2 Hyperparameter Selection for GNN
Hyperparameter search determined that 15 training iterations outperform higher values like 20 for convergence speed and final accuracy. Smaller batch sizes (32) consistently outperform larger ones, trading GPU utilization for model accuracy. With random initialization, larger batch sizes prevent 99 %+ convergence, plateauing around 90 % after 200 epochs.
Grid-structured initialization enables broader hyperparameter ranges while maintaining performance, though the final configuration uses conservative settings reliable across initialization methods.
All hyperparameters were optimized for geometric constraint problems on $\times20$ grids. A summary for GNN as well as Transformer architecture used in initial comparison reported in Section 4.1 are visible in Table 2.
Table 2: Model hyperparameters and training settings for both architectures.
Parameter GNN Transformer
Model Architecture
Embedding dimension 128 256
Model iterations/layers 15 $\pm$10 (diminishing) 6